Tuesday, 25 August

Plenary Lectures

9:00 – 10:00  Auditorium Maximum H 105

George C. Papanicolaou, Stanford University, Stanford, USA Mathematical Problems in Geophysical Wave Propagation (I: p.17)

Chair: John Ball, Oxford University, Oxford, UK

10:15 – 11:15  Auditorium Maximum H 105

Tetsuji Miwa, RIMS, Kyoto University, Kyoto, Japan Algebraic Analysis of Solvable Lattice Models (I: p.16)

Chair: Yuri Manin, Max-Planck-Institut für Mathematik, Bonn, Germany

11:45 – 12:45  Auditorium Maximum H 105

Gilles Pisier, Université Paris 6, Paris, France and Texas A&M University, College Station, USA Operator Spaces and Similarity Problems (I: p.18)

Chair: Joram Lindenstrauss, Hebrew University, Jerusalem, Israel

ICM’98 – Program  1